OpenStudy (anonymous):

Verify the identity. 4 csc 2x = 2 csc2x tan x

OpenStudy (anonymous):

@satellite73

OpenStudy (helder_edwin):

use the definitions of csc and tan.

OpenStudy (anonymous):

cscx=1/sinx

OpenStudy (helder_edwin):

\[\large \csc x=\frac{1}{\sin x}\qquad\text{and}\qquad \tan x=\frac{\sin x}{\cos x} \]

OpenStudy (helder_edwin):

and i would begin from the right side.

OpenStudy (helder_edwin):

\[\large 2\csc(2x)\tan x=2\cdot\frac{1}{\sin(2x)}\cdot\frac{\sin x}{\cos x}= \]

OpenStudy (helder_edwin):

i am sure u can finish this

OpenStudy (anonymous):

sorry its 2 csc^2x

OpenStudy (anonymous):

so would that change it to 1/sin^2x

OpenStudy (helder_edwin):

yes.

OpenStudy (anonymous):

so then would it be 2cosx2sin^2x

OpenStudy (anonymous):

@helder_edwin

OpenStudy (helder_edwin):

the identity is \[\large 4\csc(2x)=2\csc^2x\tan x \] right?

OpenStudy (anonymous):

well i meant for the right side but yes

OpenStudy (helder_edwin):

ok. so the right side would be \[\large 2\csc^2x\tan x=2\cdot\frac{1}{\sin^2x}\cdot\frac{\sin x}{\cos x} \]

OpenStudy (helder_edwin):

do u see u can simplify one sin x?

OpenStudy (anonymous):

no

OpenStudy (helder_edwin):

on the right side there is one sin(x) in the numerator and sin^2(x) in the denominator. so u can simplify one sin(x)

OpenStudy (helder_edwin):

so it becomes \[\large 2\cdot\frac{1}{\sin x}\cdot\frac{1}{\cos x}=\frac{2}{\sin x\cdot\cos x} \] agree?

OpenStudy (anonymous):

ooooooh ok

OpenStudy (helder_edwin):

now use this identity \[\large \color{red}{\sin(2x)=2\sin x\cdot\cos x} \]

OpenStudy (anonymous):

how?

OpenStudy (helder_edwin):

this can also be written as \[\large \color{red}{\frac{\sin(2x)}{2}=\sin x\cdot\cos x} \] agree?

OpenStudy (anonymous):

i guess

OpenStudy (helder_edwin):

u guess??? the 2 on the right was multiplying, it passes to the left to divide. right?

OpenStudy (helder_edwin):

anyway. if we substitute this in the identity we r trying to prove we get \[\large =\frac{2}{\sin x\cos x}=\frac{2}{\frac{\sin(2x)}{2}}=\frac{4}{\sin(2x)}= 4\cdot\csc(2x) \]

OpenStudy (anonymous):

ok yes yes

OpenStudy (helder_edwin):

got it?

OpenStudy (anonymous):

yeah i got it, can you help me with a few more?

OpenStudy (helder_edwin):

sure

OpenStudy (anonymous):

\[\frac{ 1-sinx }{ cosx }=\frac{ cosx }{ 1+sinx }\]

OpenStudy (helder_edwin):

this one is easy. say we begin from the left. ok?

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):

\[\large \frac{1-\sin x}{\cos}=\frac{1-\sin x}{\cos x}\cdot\color{red}{\frac{1+\sin x}{1+\sin x}}= \] can u finish?

OpenStudy (anonymous):

\[\frac{ 1 }{ cosx+sinx}\]??

OpenStudy (helder_edwin):

no. u have a difference of squares in the numerator, can u see it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

sorry this is like the only type of math i am bad at

OpenStudy (helder_edwin):

no problem \[\large (1-\sin x)(1+\sin x)=1^2-\sin^2x=1-\sin^2x \] agree? this is what u have in the numerator.

OpenStudy (anonymous):

oh ok, i see

OpenStudy (helder_edwin):

so we have \[\large =\frac{1-\sin^2x}{\cos(1+\sin x)}=\frac{\cos^2x}{\cos x(1+\sin x)} \] ok?

OpenStudy (anonymous):

does -sinx=cosx?

OpenStudy (helder_edwin):

no.

OpenStudy (helder_edwin):

now we can simplify one cos(x) from the numerator and one cos(x) from the denominator and get \[\large =\frac{\cos x}{1+\sin x} \] ok?

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):

cool. i hope u r really getting this stuff.

OpenStudy (anonymous):

sorta